OpenStudy (anonymous):

Help? Its urgent! Write an equation in point-slope form for the line through the given point with the given slope. (-7, 9); m=4/5

OpenStudy (anonymous):

y = mx+b form

OpenStudy (anonymous):

m represents the slope

OpenStudy (anonymous):

Use point - slope form \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

(x1,y1) are your coordinates

OpenStudy (anonymous):

y+7=4/5(x-9) y-7=4/5(x+9) y-9=4/5(x+7) y+9=4/5(x-7) what one would it be?